Write the following as numbers with base 2 :
0.25

Answer by stanbon(75887)

Write the following as numbers with base 2 :
0.25 = 1/4 = 2^-1 = 0.01
----
0.1 = 1/2 (base 2)
---
0.11 = (1/2) + 1/4 = (3/4)
---
0.01 = 1/4 = 0.25
-----------------------
Cheers,
Stan H.
===============